# COMMA: A Communicative Multimodal Multi-Agent Benchmark

### v2

@@ 5.2 Qualitative Analysis on Model Failures
Open-Source Models Have a Far Greater Tendency Towards Miscommunication
From our judge evaluations, we observe that miscommunication errors dominate the errors present in conversations produced by open-source models (60%). In comparison, GPT-4o exhibits a more balanced spread for its error distribution, with fewer miscommunication errors overall (15%). This gap in correct communication may explain the performance difference observed in Table 1 for the two models (41.74% for GPT-4o vs 22.15% for LLaMA 3.2). Many puzzles require sustained communication through various puzzle states so frequent miscommunications would lead to greater error. We hypothesize this occurs because the training data mixture for open-source models likely includes high quality single-agent data from academic benchmarks such as Visual Question Answering  (Antol et al., 2015), Image Captioning, etc. Including tasks emphasizing communication may help address this issue.
Closed-source Models Misunderstand Their Role More Often
As shown in Figure 3, the GPT-4o agent misunderstands its role in 24% of conversations compared to 9% by LLaMA 3.2. This issue can likely be addressed in future studies with more careful prompting or fine-tuning.
Both Open-source and Closed-source Models Have Similar Misinterpretation and Repetition Error Rate
A failure mode observed in both models is their tendency to repeat poor actions, often getting stuck in unproductive loops. Although GPT-4o and LLaMA 3.2 only suffer from this kind of error 6% and 4% of the time respectively, this behavior highlights a critical gap in the model’s reasoning capabilities, specifically its inability to recognize and adapt according to its conversation history. Future reasoning models may address this limitation by incorporating episodic memory during training, enabling the model to identify repeated dialogue and adapt its responses accordingly. Such a training strategy could significantly improve the agent’s ability to navigate challenging environments.
Figure 4: We plot the overall success rate on our benchmark as a function of the number of allowed conversation turns. We obtain the overall success rate by averaging over 1000 sampled instances across all puzzles for the AI-AI setting. Random is a baseline where the solver agent chooses actions uniformly at random at each time step.
@@ 5.3 Fine-grained Analysis
An important skill for agents is to learn from past mistakes to adapt to similar future situations. Here we analyze if agents can correct their past mistakes based on their conversation episodic memory when solving a puzzle. Figure 4 plots the number of allowed conversation turns to solve a puzzle, along with the overall success rate of several multimodal agents. The plot demonstrates that all models improve as conversation length increases, with top-performing models like GPT-4o, Gemini, and GPT-4V showing significantly higher improvement as conversation length increases. The open-source models such as LLaVA, InternVL, actually underperform the random baseline for most conversation lengths, suggesting limited ability to utilize episodic memory and plateau after about 4-5 conversation turns.
@@
Figure 5 evaluates the solver agent’s ability to withhold private information while successfully completing the ATM puzzle task. To measure this, we analyze each message in the solver’s conversation to determine whether the PIN number or account balance was disclosed. Our findings reveal that, despite achieving the highest success rates, both GPT-4o and Gemini frequently reveal sensitive information when interacting with the expert, even when explicitly instructed not to do so in the prompt. This suggests a significant limitation in these models’ ability to follow privacy constraints, highlighting a future growth area in enforcing strict information security in AI agent interaction.
Figure 5: Several model privacy revealing rates and success rates for the ATM puzzle across 100 runs. Despite having the best success rate, GPT-4o and Gemini also revealed the PIN most often.
Figure 6: Average partial success rate for open-source (Top) and closed-source (Bottom) models on various puzzle categories. Error bars indicate the standard error (±$\pm\sigma$) of performance for the given model on the puzzle category across 100 independent runs for each puzzle.
Performance Based on Capability
In Figure 6 we group the model performance based on the category tested: Memory (MR), Grounding (MG), Reasoning (MSR), and Reaction (PR). We observe several findings. First, open-source models perform the best on puzzles requiring grounding, and have generally worse performance on private information, memory, and multi-step reasoning puzzles. We hypothesize this is due to their lack of alignment towards these tasks during pretraining. In contrast, closed-source models excel at private information and grounding tasks, but also struggle with memory and multi-step reasoning. We again believe that despite their extensive pretraining, these models are not accustomed to using conversation history as a source of episodic memory.
@@ 6 Conclusion
In this paper, we address a critical gap in the field of multimodal agents by introducing a novel benchmark specifically designed to evaluate communication in a multimodal, multi-agent system. Our benchmark aims to simulate real-world conditions where agents possess complementary information and must work together to achieve complex goals. We comprehensively evaluate metrics such as partial success rate, mistake rate, and document common failure modes for AI-AI interactions. Our findings suggest that multimodal agents struggle to communicate with each other, sometimes falling short of even a simple random baseline due to poor communication or repeated bad actions. Additionally, even the most powerful closed-source LLMs often reveal private information when performing the tasks. These findings emphasize the need for deeper investigation into enhancing inter-agent collaboration. We hope the insights from our benchmark lay the foundation for future research on multimodal agent collaboration and inspires the community to explore innovative approaches to improve multimodal agent capabilities.
Impact Statement
While we aim to construct a holistic framework for multimodal agent communication, our experiments may not represent all possible scenarios in our puzzles. We conduct a preliminary study by sampling puzzle configurations and conversations between agents, and we leave more comprehensive evaluation of puzzle categories to future work. Additionally, there will inevitably be a simulation-to-reality gap from our benchmark to real-world situations, thus a high score on our benchmark may not perfectly generalize to real-world communication scenarios. Lastly, we acknowledge that there is inherent risk to using multimodal agents when handling private data. Given that LLMs have been shown to be prone to jailbreaking  (Chao et al., 2023; Liu et al., 2023; Shen et al., 2024), it is critical to take additional safety measures before deploying an agent in practice, even if it achieves a high benchmark score.
7 Acknowledgement
@@
Appendix A Manuals
@@
Then use the table to determine the next group to press in each stage. ”Group” refers to all squares of a particular color, or all non-white squares in the topmost row or leftmost column containing non-white squares. Pressing an incorrect square will result in a strike and reset the module. White squares will remain white for the duration of the module, but non-white squares may change color in each stage.
The table below helps to choose the next subgroup to press. The numbered keys correspond to the number of currently white squares, and the ”previously pressed color” key gives you values that indicate what color to press next based on the corresponding number of white squares.
@@
$1:$ $\ \{Blue,Column,Red,Yellow,Row,Green,Magenta\}$
$2:$ $\ \{Row,Green,Blue,Magenta,Red,Column,Yellow\}$
$3:$ $\ \{Yellow,Magenta,Green,Row,Blue,Red,Column\}$
$4:$ $\ \{Blue,Green,Yellow,Column,Red,Row,Magenta\}$
$5:$ $\ \{Yellow,Row,Blue,Magenta,Column,Red,Green\}$
$6:$ $\ \{Magenta,Red,Yellow,Green,Column,Blue,Row\}$
$7:$ $\ \{Green,Row,Column,Blue,Magenta,Yellow,Red\}$
$8:$ $\ \{Magenta,Red,Green,Blue,Yellow,Column,Row\}$
$9:$ $\ \{Column,Yellow,Red,Green,Row,Magenta,Blue\}$
$10:$ $\ \{Green,Column,Row,Red,Magenta,Blue,Yellow\}$
$11:$ $\ \{Red,Yellow,Row,Column,Green,Magenta,Blue\}$
$12:$ $\ \{Column,Row,Column,Row,Row,Column,Row\}$
$13:$ $\ \{Row,Column,Row,Column,Row,Column,Column\}$
$14:$ $\ \{Column,Column,Row,Row,Column,Row,Column\}$
$15:$ $\ \{Row,Row,Column,Row,Column,Column,Row\}$
@@
  • If the display is 1, press the button labeled ”4”.
@@
  • If the display is 4, press the button labeled ”4”.
@@
  • ”YES”: Middle Left
  • ”FIRST”: Top Right
  • ”DISPLAY”: Bottom Right
  • ”OKAY”: Top Right
  • ”SAYS”: Bottom Right
  • ”NOTHING”: Middle Left
  • ”(No Text)”: Bottom Left
  • ”BLANK”: Middle Right
  • ”NO”: Bottom Right
  • ”LED”: Middle Left
  • ”LEAD”: Bottom Right
  • ”READ”: Middle Right
  • ”RED”: Middle Right
  • ”REED”: Bottom Left
  • ”LEED”: Bottom Left
  • ”HOLD ON”: Bottom Right
  • ”YOU”: Middle Right
  • ”YOU ARE”: Bottom Right
  • ”YOUR”: Middle Right
  • ”YOU’RE”: Middle Right
  • ”UR”: Top Left
  • ”THERE”: Bottom Right
  • ”THEY’RE”: Bottom Left
  • ”THEIR”: Middle Right
  • ”THEY ARE”: Middle Left
  • ”SEE”: Bottom Right
  • ”C”: Top Right
  • ”CEE”: Bottom Right
@@
  • ”READY”: YES, OKAY, WHAT, MIDDLE, LEFT, PRESS, RIGHT, BLANK, READY, NO, FIRST, UHHH, NOTHING, WAIT
  • ”FIRST”: LEFT, OKAY, YES, MIDDLE, NO, RIGHT, NOTHING, UHHH, WAIT, READY, BLANK, WHAT, PRESS, FIRST
  • ”NO”: BLANK, UHHH, WAIT, FIRST, WHAT, READY, RIGHT, YES, NOTHING, LEFT, PRESS, OKAY, NO, MIDDLE
  • ”BLANK”: WAIT, RIGHT, OKAY, MIDDLE, BLANK, PRESS, READY, NOTHING, NO, WHAT, LEFT, UHHH, YES, FIRST
  • ”NOTHING”: UHHH, RIGHT, OKAY, MIDDLE, YES, BLANK, NO, PRESS, LEFT, WHAT, WAIT, FIRST, NOTHING, READY
  • ”YES”: OKAY, RIGHT, UHHH, MIDDLE, FIRST, WHAT, PRESS, READY, NOTHING, YES, LEFT, BLANK, NO, WAIT
  • ”WHAT”: UHHH, WHAT, LEFT, NOTHING, READY, BLANK, MIDDLE, NO, OKAY, FIRST, WAIT, YES, PRESS, RIGHT
  • ”UHHH”: READY, NOTHING, LEFT, WHAT, OKAY, YES, RIGHT, NO, PRESS, BLANK, UHHH, MIDDLE, WAIT, FIRST
  • ”LEFT”: RIGHT, LEFT, FIRST, NO, MIDDLE, YES, BLANK, WHAT, UHHH, WAIT, PRESS, READY, OKAY, NOTHING
  • ”RIGHT”: YES, NOTHING, READY, PRESS, NO, WAIT, WHAT, RIGHT, MIDDLE, LEFT, UHHH, BLANK, OKAY, FIRST
  • ”MIDDLE”: BLANK, READY, OKAY, WHAT, NOTHING, PRESS, NO, WAIT, LEFT, MIDDLE, RIGHT, FIRST, UHHH, YES
  • ”OKAY”: MIDDLE, NO, FIRST, YES, UHHH, NOTHING, WAIT, OKAY, LEFT, READY, BLANK, PRESS, WHAT, RIGHT
  • ”WAIT”: UHHH, NO, BLANK, OKAY, YES, LEFT, FIRST, PRESS, WHAT, WAIT, NOTHING, READY, RIGHT, MIDDLE
  • ”PRESS”: RIGHT, MIDDLE, YES, READY, PRESS, OKAY, NOTHING, UHHH, BLANK, LEFT, FIRST, WHAT, NO, WAIT
  • ”YOU”: SURE, YOU ARE, YOUR, YOU’RE, NEXT, UH HUH, UR, HOLD, WHAT?, YOU, UH UH, LIKE, DONE, U
  • ”YOU ARE”: YOUR, NEXT, LIKE, UH HUH, WHAT?, DONE, UH UH, HOLD, YOU, U, YOU’RE, SURE, UR, YOU ARE
  • ”YOUR”: UH UH, YOU ARE, UH HUH, YOUR, NEXT, UR, SURE, U, YOU’RE, YOU, WHAT?, HOLD, LIKE, DONE
  • ”YOU’RE”: YOU, YOU’RE, UR, NEXT, UH UH, YOU ARE, U, YOUR, WHAT?, UH HUH, SURE, DONE, LIKE, HOLD
  • ”UR”: DONE, U, UR, UH HUH, WHAT?, SURE, YOUR, HOLD, YOU’RE, LIKE, NEXT, UH UH, YOU ARE, YOU
  • ”U”: UH HUH, SURE, NEXT, WHAT?, YOU’RE, UR, UH UH, DONE, U, YOU, LIKE, HOLD, YOU ARE, YOUR
  • ”UH HUH”: UH HUH, YOUR, YOU ARE, YOU, DONE, HOLD, UH UH, NEXT, SURE, LIKE, YOU’RE, UR, U, WHAT?
  • ”UH UH”: UR, U, YOU ARE, YOU’RE, NEXT, UH UH, DONE, YOU, UH HUH, LIKE, YOUR, SURE, HOLD, WHAT?
  • ”WHAT?”: YOU, HOLD, YOU’RE, YOUR, U, DONE, UH UH, LIKE, YOU ARE, UH HUH, UR, NEXT, WHAT?, SURE
  • ”DONE”: SURE, UH HUH, NEXT, WHAT?, YOUR, UR, YOU’RE, HOLD, LIKE, YOU, U, YOU ARE, UH UH, DONE
  • ”NEXT”: WHAT?, UH HUH, UH UH, YOUR, HOLD, SURE, NEXT, LIKE, DONE, YOU ARE, UR, YOU’RE, U, YOU
  • ”HOLD”: YOU ARE, U, DONE, UH UH, YOU, UR, SURE, WHAT?, YOU’RE, NEXT, HOLD, UH HUH, YOUR, LIKE
  • ”SURE”: YOU ARE, DONE, LIKE, YOU’RE, YOU, HOLD, UH HUH, UR, SURE, U, WHAT?, NEXT, YOUR, UH UH
  • ”LIKE”: YOU’RE, NEXT, U, UR, HOLD, DONE, UH UH, WHAT?, UH HUH, YOU, LIKE, SURE, YOU ARE, YOUR

### v3

@@ 5.2 Qualitative Analysis on Model Failures
Reasoning and Open-Source Models are Worse at Communication
As shown in Figure 3, LLaMA 3.2 anad o4-mini share similar error type distributions across our benchmark, both struggling the most with communication and interpretation errors. We observed that this often occurs because the solver attempts to solve the puzzles on its own, refusing to communicate with the expert. We hypothesize this is due to the significant domain shift in pretraining data for these models, in which the data the models were trained to assume all the information is available in the prompt to answer the question. While reasoning models like o4-mini may excel at self-reflection, they still struggle with communication to perform tasks outside of the scope of their information. The open-source models display similar weaknesses in trying to solve the task on their own. To achieve a similar level of versatility as GPT-4o, adding tasks which require communication to pretraining will likely address this issue for these models.
Figure 4: Left: We plot the overall success rate on our benchmark as a function of the number of allowed conversation turns. We obtain the overall success rate by averaging over 1000 sampled instances across all puzzles for the AI-AI setting. Random is a baseline where the solver agent chooses actions uniformly at random at each time step. Right: Several model privacy revealing rates and success rates for the ATM puzzle across 100, 30, 10 runs for the AI, o4-mini, and human agents respectively. Despite having the best success rates among the AI models, GPT-4o and Gemini also revealed the PIN most often.
@@ 5.3 Fine-grained Analysis
An important skill for agents is to learn from past mistakes to adapt to similar future situations. Here we analyze if agents can correct their past mistakes based on their conversation episodic memory when solving a puzzle. Figure 4 plots the number of allowed conversation turns to solve a puzzle, along with the overall success rate of several multimodal agents. The plot demonstrates that all models improve as conversation length increases, with top-performing models like o4-mini, GPT-4o, and Gemini significantly higher improvement as conversation length increases. The open-source models such as LLaVA, InternVL, actually underperform the random baseline for most conversation lengths. Overall the results show a plateau after about 4-5 conversation turns for the AI agents, indicating that they may not use their episodic memory well. However, human agents show consistent improvement as conversation length increases.
@@
Figure 4 evaluates the solver agent’s ability to withhold private information while successfully completing the ATM puzzle task. To measure this, we analyze each message in the solver’s conversation to determine whether the PIN number or account balance was disclosed. Our analysis shows that, although models like GPT-4o and Gemini achieve relatively high success rates on the task, they frequently disclose private details to the expert agent, even when explicitly instructed to avoid doing so. In contrast, o4-mini adheres more closely to the privacy constraints, successfully withholding sensitive information, but at the cost of lower task performance. Human agents, notably, are able to solve the puzzles with perfect accuracy while maintaining complete privacy. These findings expose the inability of current multimodal models to consistently respect explicit privacy instructions and highlight a future growth area in enforcing strict information security in AI agent interaction.
Efficiency Scores and Hybrid Thinking
Figure 5 shows the tradeoff between performance and
Figure 5: Illustration of the tradeoff between performance and token usage on our benchmark. Agents in the top left corner have a better balance between performance and token usage. Among the AI agents, GPT-4o has the best balance, but still falls far below human-level efficiency.
token usage for the models we evaluated. Ideally, agents with strong communication abilities achieve high performance with minimal token usage (top left corner). We represent this property in our benchmark using the efficiency score metric, the harmonic mean between conciseness and partial score, which are the numbers in parentheses next to the model names. o4-mini displays the highest performance among all AI agents, but only has an efficiency score of 0.15 due to its long chain of thought responses. A random baseline has extremely low token usage, but poor performance, which results in an efficiency score of 0.32. Human performance shows concise and thoughtful communication, having the highest efficiency score of 0.78, which greatly surpasses the second-best score achieved by GPT-4o (0.52). We believe that incorporating hybrid thinking  Bercovich et al. (2025); Yang et al. (2025a) into these multimodal models will play an important role in improving the efficiency scores on our benchmark.
Setting Average Partial Success Rate % (↑) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Full Memory 98 ± 1.4 64 ± 4.7 72 ± 4.5 32 ± 3.4 31 ± 3.9 27 ± 2.5 2 ± 1.4 10 ± 1.6 33 ± 3.6 47 ± 5.0 41.74 ± 1.4 No Memory 95 ± 2.2 56 ± 4.8 71 ± 4.6 37 ± 3.9 24 ± 3.1 24 ± 2.3 7 ± 2.6 8 ± 1.3 3 ± 1.7 45 ± 5.5 36.20 ± 1.4 Single Agent 82 ± 3.9 74 ± 4.3 56 ± 5.0 66 ± 3.6 24 ± 3.0 34 ± 3.7 11 ± 3.1 17 ± 1.6 33 ± 5.0 8 ± 2.7 40.41 ± 1.4
Table 2: Impact of changing information access on performance in our benchmark for a GPT-4o agent. Interestingly, providing the agent with full information slightly worsens the overall performance. However, performance on more reasoning-heavy tasks improves. Withholding memory from past conversations hurts performance across most puzzles.
Effect of Memory and View Asymmetry
To isolate the effects of memory input and view asymmetry, we run two ablations on the collaborative setting. First, we remove episodic memory by restricting the agent to only the latest conversation message. Second, we remove view asymmetry by giving a single agent full access to both the solution manual and the image. Results are shown in Table 2. Removing memory significantly reduces performance, especially on tasks requiring reasoning across multiple turns (e.g., maze and memory puzzles). This shows that collective memory is crucial for recalling information over interactions; without it, the agent repeats itself, contradicts earlier steps, or loses progress. Interestingly, the single-agent setting with full information also performed slightly worse overall, though it sometimes outperformed collaboration on reasoning-heavy tasks (e.g., LED, Color). The large performance gap across tasks suggests that while some benefit from collaboration, others are better handled by individual reasoning.
@@ 6 Conclusion
In this paper, we address a critical gap in the field of multimodal agents by introducing a novel benchmark specifically designed to evaluate communication in a multimodal, multi-agent system. Our benchmark aims to simulate real-world conditions in which agents possess complementary information and must work together to achieve complex goals. We comprehensively evaluate metrics such as partial success rate and efficiency score, and document common failure modes for AI-AI interactions. Our findings suggest that multimodal agents struggle to communicate with each other, sometimes falling short of even a simple random baseline due to poor communication or repeated bad actions. This issue is especially apparent in recent reasoning models which tend to ignore other agents and try to solve the puzzles by themselves. Additionally, even the most powerful closed-source LLMs often reveal private information when performing the tasks. These findings emphasize the need for deeper investigation into enhancing inter-agent collaboration. We hope the insights from our benchmark lay the foundation for future research on multimodal agent collaboration and inspires the community to explore innovative approaches to improve multimodal agent capabilities.
7 Broader Impact and Limitations
Privacy and Sensitive Data Risks
Our benchmark includes puzzles involving sensitive user data, such as health and financial information (e.g., TelehealthPuzzle, ATMPuzzle). Although these scenarios are simulated, they are designed to probe agent behavior in privacy-critical contexts. We observed that even high-performing models, such as GPT-4o and Gemini 2.0, frequently fail to follow privacy-preserving instructions, disclosing confidential information despite explicit constraints in their prompt. This suggests an ethical concern where current models may not be suitable for deployment in domains where privacy is essential. Additionally, strong benchmark performance does not equate to safe behavior in real-world settings, and developers should implement robust privacy-preserving mechanisms before considering deployment.
Simulation-to-Reality Gap
Similar to other benchmarks, our framework necessarily simplifies real-world complexity. While our experiments involve sampling puzzle configurations and simulating multi-turn agent conversations, they do not exhaust the vast space of all possible collaborative interactions. For instance, real-world scenarios often involve collaboration with more than 2 people and hierarchical role structures in corporate environments. To help bridge this gap, future work may consider adding more realistic collaboration environments. First, implement a hierarchical structure of agent communication, such as a workplace of multiple agents for developing software. Second, incorporating real-world modality data (e.g., images, audio, or video) may help simulate a more faithful multimodal setting. For example, interacting with medical imaging software would be much more complicated than our current puzzles, but simulate a realistic scenario. Lastly, the inclusion of more human-in-the-loop evaluation of performance can provide more accurate insight.
LLM Judge Evaluation Bias
While our main experiments utilize rule-based evaluation to calculate success rate statistics, we use o1 as an automatic judge to classify conversation error types, including conversations from models in the same architectural family. This introduces the risk of unintended evaluation bias. Although automated judging enables scalable analysis, it may not be fully impartial.
Human Baseline and Author Bias
The human baseline in our study was established using a small sample consisting of the authors who also designed the tasks. This creates the possibility of unconscious bias and limits the generalizability of our human-agent comparisons. While this setup was sufficient for initial prototyping, we recognize that it may over or underestimate the performance gap between humans and agents. Despite these constraints, we believe that our current human evaluation still provides a useful signal, which is an upper bound of human performance under idealized conditions. While average performance from a broader pool of external participants may be lower, our results nevertheless reveal consistent trends in human responses. First, our current human baseline reflects a tendency for human communication to be more concise than that of language models, a pattern likely to persist even after the inclusion of more human solver trials. Second, our human participants show nearly perfect vision perception and describe image content with high precision. To improve this human baseline, several aspects can be considered in future work. First, the codebase can be redesigned for web-based deployment and crowdsourcing a diverse pool of participants. Second, in addition to random puzzle initialization, the implementation of randomized rules in the instruction manual may help curate unbiased human performance metrics.
Responsible Use and Deployment
Agents that perform well on benchmarks may still behave unpredictably when deployed, especially in sensitive domains like healthcare. Language-only interfaces, while convenient for simulation, can obscure the nuances of real-world decision-making, particularly when human lives or well-being are involved. We strongly caution against deploying collaborative agents in such contexts without extensive experimental trials.
Acknowledgement
@@
Appendix A Puzzle Details
Example Dialog and Solver-Expert Interaction Details
We include a full example trajectory between the solver and expert on a memory puzzle in Figure 6. At each turn, the solver receives a multimodal input by concatenating the instruction, dialogue history, and image tokens from the puzzle and generates a textual response representing its next action. The expert then takes the updated context, including the solver’s response, and replies in a similar manner. This turn-based exchange continues throughout the episode. Our framework manages this interaction without assuming any shared architecture between agents, allowing flexible use of different vision-language models.
Figure 6: A full trajectory example of a Gemini solver (green) and expert (red) on an instance of the memory puzzle. The solver has access to the image at the current puzzle state at each time step, while the expert has access to the solution manual at all times. At each dialogue step, each agent also receives the entire conversation history as part of its input.
@@ Appendix A Puzzle Details
Then use the table to determine the next group to press in each stage. "Group" refers to all squares of a particular color, or all non-white squares in the topmost row or leftmost column containing non-white squares. Pressing an incorrect square will result in a strike and reset the module. White squares will remain white for the duration of the module, but non-white squares may change color in each stage.
The table below helps to choose the next subgroup to press. The numbered keys correspond to the number of currently white squares, and the "previously pressed color" key gives you values that indicate what color to press next based on the corresponding number of white squares.
@@
  • If the display is 1, press the button labeled "4".
@@
  • If the display is 4, press the button labeled "4".
@@
  • "YES": Middle Left
  • "FIRST": Top Right
  • "DISPLAY": Bottom Right
  • "OKAY": Top Right
  • "SAYS": Bottom Right
  • "NOTHING": Middle Left
  • "(No Text)": Bottom Left
  • "BLANK": Middle Right
  • "NO": Bottom Right
  • "LED": Middle Left
  • "LEAD": Bottom Right
  • "READ": Middle Right
  • "RED": Middle Right
  • "REED": Bottom Left
  • "LEED": Bottom Left
  • "HOLD ON": Bottom Right
  • "YOU": Middle Right
  • "YOU ARE": Bottom Right
  • "YOUR": Middle Right
  • "YOU’RE": Middle Right
  • "UR": Top Left
  • "THERE": Bottom Right
  • "THEY’RE": Bottom Left
  • "THEIR": Middle Right
  • "THEY ARE": Middle Left
  • "SEE": Bottom Right
  • "C": Top Right
  • "CEE": Bottom Right
@@
  • "READY": YES, OKAY, WHAT, MIDDLE, LEFT, PRESS, RIGHT, BLANK, READY, NO, FIRST, UHHH, NOTHING, WAIT
  • "FIRST": LEFT, OKAY, YES, MIDDLE, NO, RIGHT, NOTHING, UHHH, WAIT, READY, BLANK, WHAT, PRESS, FIRST
  • "NO": BLANK, UHHH, WAIT, FIRST, WHAT, READY, RIGHT, YES, NOTHING, LEFT, PRESS, OKAY, NO, MIDDLE
  • "BLANK": WAIT, RIGHT, OKAY, MIDDLE, BLANK, PRESS, READY, NOTHING, NO, WHAT, LEFT, UHHH, YES, FIRST
  • "NOTHING": UHHH, RIGHT, OKAY, MIDDLE, YES, BLANK, NO, PRESS, LEFT, WHAT, WAIT, FIRST, NOTHING, READY
  • "YES": OKAY, RIGHT, UHHH, MIDDLE, FIRST, WHAT, PRESS, READY, NOTHING, YES, LEFT, BLANK, NO, WAIT
  • "WHAT": UHHH, WHAT, LEFT, NOTHING, READY, BLANK, MIDDLE, NO, OKAY, FIRST, WAIT, YES, PRESS, RIGHT
  • "UHHH": READY, NOTHING, LEFT, WHAT, OKAY, YES, RIGHT, NO, PRESS, BLANK, UHHH, MIDDLE, WAIT, FIRST
  • "LEFT": RIGHT, LEFT, FIRST, NO, MIDDLE, YES, BLANK, WHAT, UHHH, WAIT, PRESS, READY, OKAY, NOTHING
  • "RIGHT": YES, NOTHING, READY, PRESS, NO, WAIT, WHAT, RIGHT, MIDDLE, LEFT, UHHH, BLANK, OKAY, FIRST
  • "MIDDLE": BLANK, READY, OKAY, WHAT, NOTHING, PRESS, NO, WAIT, LEFT, MIDDLE, RIGHT, FIRST, UHHH, YES
  • "OKAY": MIDDLE, NO, FIRST, YES, UHHH, NOTHING, WAIT, OKAY, LEFT, READY, BLANK, PRESS, WHAT, RIGHT
  • "WAIT": UHHH, NO, BLANK, OKAY, YES, LEFT, FIRST, PRESS, WHAT, WAIT, NOTHING, READY, RIGHT, MIDDLE
  • "PRESS": RIGHT, MIDDLE, YES, READY, PRESS, OKAY, NOTHING, UHHH, BLANK, LEFT, FIRST, WHAT, NO, WAIT
  • "YOU": SURE, YOU ARE, YOUR, YOU’RE, NEXT, UH HUH, UR, HOLD, WHAT?, YOU, UH UH, LIKE, DONE, U
  • "YOU ARE": YOUR, NEXT, LIKE, UH HUH, WHAT?, DONE, UH UH, HOLD, YOU, U, YOU’RE, SURE, UR, YOU ARE
  • "YOUR": UH UH, YOU ARE, UH HUH, YOUR, NEXT, UR, SURE, U, YOU’RE, YOU, WHAT?, HOLD, LIKE, DONE
  • "YOU’RE": YOU, YOU’RE, UR, NEXT, UH UH, YOU ARE, U, YOUR, WHAT?, UH HUH, SURE, DONE, LIKE, HOLD
  • "UR": DONE, U, UR, UH HUH, WHAT?, SURE, YOUR, HOLD, YOU’RE, LIKE, NEXT, UH UH, YOU ARE, YOU
  • "U": UH HUH, SURE, NEXT, WHAT?, YOU’RE, UR, UH UH, DONE, U, YOU, LIKE, HOLD, YOU ARE, YOUR
  • "UH HUH": UH HUH, YOUR, YOU ARE, YOU, DONE, HOLD, UH UH, NEXT, SURE, LIKE, YOU’RE, UR, U, WHAT?
  • "UH UH": UR, U, YOU ARE, YOU’RE, NEXT, UH UH, DONE, YOU, UH HUH, LIKE, YOUR, SURE, HOLD, WHAT?
  • "WHAT?": YOU, HOLD, YOU’RE, YOUR, U, DONE, UH UH, LIKE, YOU ARE, UH HUH, UR, NEXT, WHAT?, SURE
  • "DONE": SURE, UH HUH, NEXT, WHAT?, YOUR, UR, YOU’RE, HOLD, LIKE, YOU, U, YOU ARE, UH UH, DONE
  • "NEXT": WHAT?, UH HUH, UH UH, YOUR, HOLD, SURE, NEXT, LIKE, DONE, YOU ARE, UR, YOU’RE, U, YOU
  • "HOLD": YOU ARE, U, DONE, UH UH, YOU, UR, SURE, WHAT?, YOU’RE, NEXT, HOLD, UH HUH, YOUR, LIKE
  • "SURE": YOU ARE, DONE, LIKE, YOU’RE, YOU, HOLD, UH HUH, UR, SURE, U, WHAT?, NEXT, YOUR, UH UH
  • "LIKE": YOU’RE, NEXT, U, UR, HOLD, DONE, UH UH, WHAT?, UH HUH, YOU, LIKE, SURE, YOU ARE, YOUR